OpenStudy (anonymous):

To solve for w in the formula V = lwh, you multiply both sides by lh. true or false

OpenStudy (anonymous):

\[ V = l*w*h\]\[ \frac{V}{l*h} = \frac{l*w*h}{l*h}\]\[ \frac{V}{l*h} = w\]

OpenStudy (anonymous):

what did I do there?

OpenStudy (unklerhaukus):

divide instead of multiply, like romero has done